# Heterogeneous Multi-Agent Reinforcement Learning for Zero-Shot Scalable Collaboration

### v1

@@ IV-B Heterogeneous Layer Design
With the generated latent variable for each agent, we can further design the heterogeneous layer for each agent described in Fig. 2(c, d). The heterogeneous layer is a linear layer in the ActorNet, which also has an MLP and an RNN ahead and an MLP behind. The ActorNet for agent $n$ takes the observation $o_{n}$ and the hidden states of the RNN $h^{t-1}_{n}$ to get the action $a_{n}$. The ActorNet for different agents shares the parameters, except for the heterogeneous layer, whose parameters are generated with each agent’s $l$.
The heterogeneous layer computes $hete\_f_{n}$ as Equation 5.
$hete\_f_{n}=w_{n}*o\_f_{n}+b_{n}.$ (5)
Here, the weights $w_{n}$ and the bias $b_{n}$ are unique for different agents, so the decision-making policies are heterogeneous for them. With a decoder implemented as an MLP and two different linear decoders to generate $w_{n}$ and $b_{n}$ based on $l_{n}$, respectively. Namely, the LatentNet learns to represent the strategy patterns, and the heterogeneous layer turns the latent variable into parameters of the ActorNet. Now, the agents can have distinguished styles to take action, though they share most of the parameters. Also, they can change the style adaptively as the task goes on or team size changes. The execution of one time step is shown in Algorithm 1.
@@ IV-C Loss Formulation
A good latent representation should be identifiable, diverse, and helpful for decision-making. Therefore, we design three distinct loss items to guide the learning of the LatentNet.
For the first two demands, we apply unsupervised losses to learn the latent variable. The first item is $\mathcal{L}_{e}$, the mean entropy of the $\boldsymbol{\mathcal{D}}_{i}$, defined as follows:
@@
The third item is the value $V_{I}$ from the InferenceNet $\mathcal{L}_{v}$:
@@
$\mathcal{L}_{L}(\theta_{L})=\lambda_{e}\mathcal{L}_{e}(\theta_{L})-\lambda_{d}\mathcal{L}_{d}(\theta_{L})-\lambda_{v}\mathcal{L}_{v}(\theta_{L}),$ (10)
where $\lambda_{e},\lambda_{d}$, and $\lambda_{v}$ are the weights of the three items of the final loss.
The InferenceNet is learned by an MSE loss between the value $V_{I}$ and environment return $R$:
$\mathcal{L}_{I}(\theta_{I})=$ $MSE\_Loss[V_{I}(o_{g},\boldsymbol{\mu}(\boldsymbol{o},\boldsymbol{h}^{t-1};\theta_{I}),$
$\boldsymbol{\sigma}(\boldsymbol{o},\boldsymbol{h}^{t-1};\theta_{I})),R],$ (11)
@@
note that here $\boldsymbol{l}$ is taken as an input without gradients.
@@
IV-D Overall Training Procedure
In one episode, the first phase is the interaction with the environment to collect data. Every agent gains a latent variable and then decides its action decision at each step (see Algorithm 1). Data such as observations, hidden states, latent variables, actions, and rewards are saved in the buffer. After the agents succeed, fail, or the step limit is reached, the second phase is training (see Algorithm 2). The agents update their ActorNets one by one, following the procedure in HAPPO [14]. Of course, here all the agents share the same ActorNet. Thus, the ActorNet will be updated for $n$ times based on different agents’ observations, actions, and rewards. After that, the CriticNet, the LatentNet, and the InferenceNet update their parameters. Please refer to Appendix  B and C for the detailed network and hyperparameter configurations.
@@ IV-D Overall Training Procedure
        Push transitions $\{(o^{t}_{i},h^{t-1}_{i},l^{t}_{i},a^{t}_{i},o^{t+1}_{i},r^{t})\},0\leq t\leq T,i=1,...,n$ into $\mathcal{B}$
        Sample a random minibatch of $B$ transitions from $\mathcal{B}$
@@
        Update CriticNet as Equation IV-C
@@
            $\mathcal{L}_{L}(\theta_{L})=\lambda_{e}\mathcal{L}_{e}(\theta_{L})-\lambda_{d}\mathcal{L}_{d}(\theta_{L})-\lambda_{v}\mathcal{L}_{v}(\theta_{L})$
        Update InferenceNet by $\mathcal{L}_{I}(\theta_{I})$ as Equation IV-C
@@ V-A Environments and Metrics
Google Research Football (GRF) contains diverse subtasks in a football game. The agents cooperate as part of a football team to score a goal against the opposing team under different scenarios such as counterattack, shooting in the penalty area, etc.
@@
Task Metric SHPPO HAPPO HAPPO (share)* MAPPO (share) HATRPO (share)
@@
In order to directly zero-shot transfer the learned models to tasks with varied populations, we modify the original environments to fix the number of allies and enemies that one agent can observe. This way, the observations and actions length is the same before and after the transfer. Hence, we can concentrate on transferring strategies without the need to map observations, which is not the primary focus of this paper. For example, in the task MMM2 of SAMC, there are 10 controlled agents and 12 enemies, respectively. In the original tasks of SMAC, the agent can observe all the enemies, resulting in easy and unscalable tasks. We modify the tasks to limit the observation range to the closest 10 enemies and the closest 8 allies so that we can create a series of new tasks with different numbers of agents and enemies to test scalability. While in GRF, there are always 11 players in one team. We change the number of players that the algorithm can control to test scalability but keep the observations as the global observations of all the players in the field. More details of the modified tasks can be found in Appendix Table IV. All the experiments in our paper are conducted on the new scalable tasks. As for metrics, we adopt training rewards to evaluate the training progress. We also test the performance on the test environments to calculate the win rate for SMAC or score rate for GRF.
We compare the performance of our approach with four other baselines. As SHPPO is built based on parameter-shared HAPPO, we take both HAPPO and HAPPO (share) as baselines. HATRPO [14] is another heterogeneous MARL method, and for a fair comparison, we adopt its parameter-shared version as a baseline. MAPPO [20] is a classic homogeneous MARL method, and in this paper, we also use parameter-shared MAPPO.
V-B Results on SMAC
We first evaluate SHPPO on the original tasks. The first task is a heterogeneous task MMM2, where each side has three kinds of different units - Marine, Marauder, and Medivac (Fig. 3(a)). They have different health points, hit points, and even special skills. Therefore, heterogeneous strategies will improve the team’s performance by maximizing each agent’s advantages. In Fig. 4(a), HAPPO performs best for its great heterogeneous representation ability. However, HAPPO models every agent individually and thus lacks scalability. In contrast, our approach, SHPPO, outperforms all the other parameter-shared baselines without modeling each agent. The training reward and win rate of SHPPO are very close to those of HAPPO, showing that SHPPO has a similar heterogeneous representation ability to HAPPO (See Table I). The performance of the two heterogeneous methods, HAPPO and HATRPO, is the worst when parameters are shared.
We also evaluate on a homogeneous task 8m_vs_9m (Fig. 3(b)), where each side has the same kind of unit - Marine. Though the agents are all the same, they can have different strategies and adaptively alter during the task. As shown in Fig. 4(b) and Table I, SHPPO outperforms all the baselines including HAPPO. The training process of SHPPO also converges faster than other baselines.
Figure 4: Performance on SMAC and GRF. We plot the win rate on SMAC and the score rate on GRF during training. (a) MMM2 (SMAC). (b) 8m_vs_9m (SMAC). (c) 3_vs_1_with_keeper (GRF). (d) counterattack_easy (GRF). The confidence interval is calculated on 5 seeds.
TABLE II: Scalability test on SMAC.
Task* SHPPO HAPPO** HAPPO (share) MAPPO (share) HATRPO (share)
@@ V-B Results on SMAC
    When transferring HAPPO, We randomly select a model of the same type for the added agent, or removing extra models when there are fewer units in the new task.
@@ V-C Results on GRF
In another task counterattack_easy (Fig. 3(d)), four agents need to cooperate together to do a counterattack. They need to pass the ball against one adversary player and the keeper to score. In Fig. 4(d), HAPPO converges fastest and has the best performance. SHPPO reaches a similar performance, though at a slow speed. Still, SHPPO outperforms all the parameter-shared baselines (also see Table I).
Figure 5: Ablation studies. The experiments are conducted on MMM2 and 8m_vs_9m in SMAC. (a, b) Ablation studies of LatentNet’s inputs. (c, d) Ablation studies of LatentNet’s losses. The confidence interval is calculated on 5 seeds.
TABLE III: Scalability test on GRF.
Task SHPPO HAPPO HAPPO (share) MAPPO (share) HATRPO (share)
@@ V-D Sacability Test
From Table II, SHPPO performs the best on most of the unseen zero-shot tasks. Even in cases where SHPPO is not optimal, it still performs comparably to the best. It’s worth noting that although we attempt to transfer HAPPO, it only outperforms SHPPO in one scenario (711_821), despite having a better win rate on the original task.
@@
Similarly, we test scalability by removing or adding one agent in the two tasks of GRF. SHPPO maintains the best performance after transfer on the 3_vs_1_with_keeper series. For the counterattack_easy series, despite HAPPO’s best performance on the original task, SHPPO achieves the superior score rate on 3_vs_1_counterattack_with_keeper after the transfer. And when adding a new agent, SHPPO also has comparable performance.
@@ V-E Ablation Studies
In order to test the effectiveness of the latent variables, we first only mask the inputs of the LatentNet with zeros so that the latent variables cannot be adaptively updated according to the new observations. Moreover, in another experiment, we replace all the latent variables with zeros to totally remove the latent learning. In Fig. 5(a, b), when the inputs of LatentNets are zeros, the performance on two tasks of SMAC reduces significantly, especially in task 8m_vs_9m. When the latent variables are set as zeros, it is equal to the removal of the latent learning part, and we can see even worse performance in such a case. Therefore, heterogeneous strategies and flexible adjustment of strategies are important for the agents to cooperate.
@@
Next, we test the effectiveness of the losses of latent learning. We first remove the guidance from the InferenceNet $\mathcal{L}_{v}(\theta_{L})$ while keeping the unsupervised losses $\mathcal{L}_{e}(\theta_{L})$ and $\mathcal{L}_{d}(\theta_{L})$. Results in Fig. 5(c, d) show that without the InferenceNet, there will be a performance decline on both MMM2 and 8m_vs_9m. It suggests that the InferenceNet is able to assess the value of latent variables to help with agents’ decision-making.
@@
Similarly, in Fig. 5(c, d), we then keep $\mathcal{L}_{v}(\theta_{L})$ but remove $\mathcal{L}_{e}(\theta_{L})$ and $\mathcal{L}_{d}(\theta_{L})$, respectively. The two experiments have similar performance reductions on task MMM2. Nevertheless, on the homogeneous task 8m_vs_9m, the experiment without the entropy loss $\mathcal{L}_{e}(\theta_{L})$ has the lowest win rate. We believe this is because, the strategies for a homogeneous task may have some extent of diversity for labor division, but the strategies cannot be too random and different as the agents are of the same unit type. While the entropy loss $\mathcal{L}_{e}(\theta_{L})$ can make sure that the learned latent variables are relatively deterministic and identifiable.
V-F Strategy Analysis
Figure 6: Latent variables in one episode in the original and unseen tasks with different population sizes. The latent variables have three dimensions X, Y, Z and change dynamically with the steps. All the ally agents are illustrated and we use colors to distinguish the unit type. The new agent is marked dots. (a) The original task MMM2. (b) The transferred task 722_831 with a new Medivac. (c) The transferred task 821_831 with a new Marine. (d) The transferred task 731_831 with a new Marauder.
Figure 7: Visualization of MMM2 and latent variables. The agents adaptively adopt different strategies during the task. The latent variables cluster accordingly. Red: stay back to be covered. Green: attack in the distance. Blue: attract the fire by moving to the front.
Figure 8: Visualization of the transferred task 821_831 and latent variables. The agents adaptively adjust strategies after being transferred to an unseen task with one added Marine (Number 9). The latent variables cluster accordingly. Green: attack in the distance. Blue: attract the fire by moving to the front.
Finally, we visualize the learned latent variables and the progress of the task to analyze the heterogeneous strategies during the task. Take MMM2 as an example. In Fig. 6, we plot how the latent variables update in one episode. In the original task (Fig. 6(a)), the latent variable patterns vary among different types of units. The two Marauders’ latent variables basically have the same trends, while those of the Marines are much more diverse. Most of the agents do not stick to only one strategy during the task. Instead, they adaptively alter the latent variables to change the strategy. With the learned heterogeneous strategies, we transfer the models to three unseen tasks adding one agent of different types (Fig. 6(b, c, d)). We can see that, generally, the agents’ latent variables still follow similar heterogeneous patterns to those in the original task, with slight differences in strategy updates during the task. Thanks to the shared LatentNet, the newly added agent can initially adopt one of the strategies from units of the same type. However, over time, the new agents may also develop novel strategies to enhance their performance in the new task by temporal heterogeneity (Fig. 6(b) after step 20, Fig. 6(c) after step 15).
We also illustrate the screenshots with the corresponding latent variables before (Fig. 7) and after (Fig. 8) the transfer. In Fig. 7, at the early stage of the task (Step 5), the agents form two groups, the Marines move forward and attack in the front (green circle) to cover the Marauders and the Medivac behind (red circle). At step 11, the two armies encounter and the strategies alter accordingly. The Marauders get close to the enemies to attract the fire while the Medivac keeps curing them (blue circle). Some Marines (Agent 2, 5, and 8) are hurt with low health points, thus they retreat to be covered (red circle), while the other Marines attack the enemies (green circle). In the final stage, some agents are dead and the enemies are almost defeated. The only Marauder keeps attracting the fire in the front. However, the Marines attack the enemies regardless of their health points to achieve final victory.
In Fig. 8, one new Marine is added while the enemies are the same, making the new task easier. Thus, the agents inherit the previous strategies and roles in the team, but further adaptively align them with the easier scenario. At step 7, the added Marine (Agent 9) joins its fellow Marines to attack the enemies (green circle), while other Marines are moving forward to attract and disperse the enemies’ fires, together with the Medivac and Marauders (blue circle). Note that no agents keep staying back to be covered as the red circle in Fig. 7, because they are taking a more aggressive strategy in easier tasks. At step 16, all the Marines stop moving and focus on attacks (green circle). The Marauders go on moving to the front with the support of the Medivac (blue circle). At the final stage (step 33), Marauder 1 dies, and almost at the same time, Marine 3 changes its strategy to replace the dead Marauder, moving to the front (blue circle).
We can see that the latent variables successfully learn the heterogeneous strategy patterns in this task, which are located diversely in the latent space. These latent variables contribute to the adaptive policy updates during the task through the HeteLayer in the ActorNet, so the agents can not only have distinct strategies but also change their strategies as the task progresses and the population size varies.

### v2

@@ IV-B Heterogeneous Layer Design
With the generated latent variable for each agent, we can further design the heterogeneous layer for each agent described in Fig. 2(c, d). The heterogeneous layer is a linear layer in the ActorNet. The ActorNet for agent $i$ takes the observation $o_{i}$ and the hidden states of the RNN $h^{t-1}_{i}$ to get the action $a_{i}$. Here, in order to make better sequential decisions, the RNN is introduced to keep the memory from the previous steps by the hidden states. The ActorNet for different agents shares the parameters, except for the heterogeneous layer, whose parameters are generated with each agent’s $l$. The heterogeneous layer computes $hete\_f_{i}$ as Equation 5.
$hete\_f_{i}=w_{i}*o\_f_{i}+b_{i}.$ (5)
Here, the weights $w_{i}$ and the bias $b_{i}$ are distinct for each agent, resulting in heterogeneous decision-making policies. A decoder, implemented as an MLP, along with two separate linear decoders (denoted as w_decoder and b_decoder), generates $w_{i}$ and $b_{i}$ based on $l_{i}$. In sum, the LatentNet learns to represent the strategy patterns, and the heterogeneous layer turns the latent variable into part of the ActorNet’s parameters. This allows agents to exhibit unique action styles, despite sharing the majority of parameters. Furthermore, they can adapt their styles in response to evolving tasks or changes in team size. The execution of one time step is shown in Algorithm 1.
@@ IV-C Loss Formulation
In this subsection, we formulate the losses to train the networks deployed in our framework. A good latent representation should be helpful for decision-making, and also identifiable and diverse. Therefore, we design three distinct loss items to guide the learning of the LatentNet.
We first employ the value $V_{I}$ from the InferenceNet $\mathcal{L}_{v}$ to evaluate the impacts of the latent variables on decision-making:
Then, we incorporate two additional unsupervised losses as regularization terms to facilitate the learning of identifiable and diverse latent variables.
The first item is $\mathcal{L}_{e}$, the mean entropy of the $\boldsymbol{\mathcal{D}}_{i}$, defined as follows:
@@
$\mathcal{L}_{L}(\theta_{L})=-\mathcal{L}_{v}(\theta_{L})+\lambda_{e}\mathcal{L}_{e}(\theta_{L})-\lambda_{d}\mathcal{L}_{d}(\theta_{L}),$ (10)
where $\lambda_{e}$ and $\lambda_{d}$ are the weights of the regularization items. Please see Appendix Table V for the details of the weights.
The InferenceNet is learned by a mean squared error (MSE) loss between the value $V_{I}$ and environment return $R$:
$\mathcal{L}_{I}(\theta_{I})=$ $MSE\_Loss[V_{I}(o_{g},\boldsymbol{\mu},\boldsymbol{\sigma};\theta_{I}),R],$ (11)
@@
note that here $\boldsymbol{l}$ is taken as an input without gradients. $V_{C}$ is the value from the critic.
@@
IV-D Efficacy of the LatentNet and InferenceNet
In this subsection, we further explain how the latent learning works to adaptively represent the agents’ heterogeneous strategies in a low-dimension space as the latent variables.
IV-D1 InferenceNet
The InferenceNet is trained to approximate a value function $V_{I}$ based on the latent representations of all the agents $\boldsymbol{\mu}(\boldsymbol{o},\boldsymbol{h}^{t-1};\theta_{L})$ and $\boldsymbol{\sigma}(\boldsymbol{o},\boldsymbol{h}^{t-1};\theta_{L})$ produced by the LatentNet, as well as the global observation $o_{g}$. The objective is to minimize the difference between the predicted value $V_{I}$ and the true return $R$. This difference is captured by the MSE loss function (Equation 11).
The gradient of this loss function $\mathcal{L}_{I}$ with respect to the parameters $\theta_{I}$ of the InferenceNet is:
$\nabla_{\theta_{I}}\mathcal{L}_{I}(\theta_{I})$ $=2\mathbb{E}[(V_{I}-R)\nabla_{\theta_{I}}V_{I}].$ (15)
By iteratively applying gradient descent, the parameters $\theta_{I}$ are adjusted such that the value function $V_{I}$ better approximates the true return $R$. Namely, the value $V_{I}$ centrally assesses the overall adequacy of the learned latent distributions of the entire team in facilitating decision-making processes based on global observations.
IV-D2 LatentNet
With the value $V_{I}$ as one of the optimization targets, the LatentNet learns better latent representations to maximize the value as Equation 10. The gradient of this loss function $\mathcal{L}_{v}$ with respect to the parameters $\theta_{L}$ of the LatentNet is given by:
$\nabla_{\theta_{L}}\mathcal{L}_{v}(\theta_{L})$ $=\nabla_{\theta_{L}}V_{I}(o_{g},\mu(o,h_{t-1};\theta_{L}),\sigma(o,h_{t-1};\theta_{L}))$
$=\frac{\partial V_{I}}{\partial\mu}\cdot\frac{\partial\mu}{\partial\theta_{L}}+\frac{\partial V_{I}}{\partial\sigma}\cdot\frac{\partial\sigma}{\partial\theta_{L}}.$ (16)
Here, $\frac{\partial V_{I}}{\partial\mu}$ and $\frac{\partial V_{I}}{\partial\sigma}$ represent the sensitivity of the value function $V_{I}$ to the changes in the latent representations $\mu$ and $\sigma$, respectively. Meanwhile, $\frac{\partial\mu}{\partial\theta_{L}}$ and $\frac{\partial\sigma}{\partial\theta_{L}}$ denote the dependence of these latent representations on the parameters of the LatentNet. By maximizing this loss, the LatentNet learns to produce latent vectors $\mu$ and $\sigma$ that are more effective in representing the policy, thereby enhancing the overall decision-making process.
In conclusion, the output of the InferenceNet, which serves as the loss for the LatentNet, establishes a direct feedback mechanism that aligns the latent representations with the objective of maximizing the expected value function, resulting in enhanced policy performance. Additionally, when combined with the two other unsupervised losses, $\mathcal{L}_{e}$ and $\mathcal{L}_{d}$, the latent representations become more identifiable and diverse.
IV-E Overall Training Procedure
In one episode, the first phase is the interaction with the environment to collect data. Every agent gains a latent variable and then decides its action decision at each step (see Algorithm 1). Data such as observations, hidden states, latent variables, actions, and rewards are saved in the buffer. After the agents succeed, fail, or the step limit is reached, the second phase is training (see Algorithm 2). The agents update their ActorNets one by one, following the procedure in HAPPO [14]. Of course, here all the agents share the same ActorNet. Thus, the ActorNet will be updated for $n$ times based on different agents’ observations, actions, and rewards. After that, the CriticNet, the LatentNet, and the InferenceNet update their parameters. Please refer to Appendix Table IV and V for the detailed network and hyperparameter configurations.
@@ IV-E Overall Training Procedure
        Push transitions $\{(o^{t}_{i},h^{t-1}_{i},l^{t}_{i},a^{t}_{i},o^{t+1}_{i},r^{t})\},0\leq t\leq T,i=1,...,n$ into $\mathcal{B}$.
        Sample a random minibatch of $B$ transitions from $\mathcal{B}$.
@@
        Update CriticNet as Equation IV-C.
        Compute $\mathcal{L}_{v}(\theta_{L})$ by InferenceNet as Equation 6.
@@
            $\mathcal{L}_{L}(\theta_{L})=-\mathcal{L}_{v}(\theta_{L})+\lambda_{e}\mathcal{L}_{e}(\theta_{L})-\lambda_{d}\mathcal{L}_{d}(\theta_{L})$.
        Update InferenceNet by $\mathcal{L}_{I}(\theta_{I})$ as Equation 11.
@@ V-A Environments and Metrics
Google Research Football (GRF) contains diverse subtasks in a football game. The agents cooperate as part of a football team to score a goal against the opposing team under scenarios such as counterattack, shooting in the penalty area, etc.
Figure 4: Performance on SMAC and GRF. We plot the win rate on SMAC and the score rate on GRF during training. (a) MMM2 (SMAC). (b) 8m_vs_9m (SMAC). (c) 3_vs_1_with_keeper (GRF). (d) counterattack_easy (GRF). The confidence interval is calculated over 5 seeds.
@@
Task Metric SHPPO (ours) HAPPO HAPPO (share)* MAPPO (share) HATRPO (share)
@@
In order to directly zero-shot transfer the learned models to tasks with varied populations, we modify the original environments to fix the number of allies and enemies that one agent can observe. This way, the observations and actions length is the same before and after the transfer. Hence, we can concentrate on transferring strategies without the need to map observations, which is not the primary focus of this paper. More details of the modified tasks can be found in Appendix A. All the experiments in our paper are conducted on the new scalable tasks.
As for metrics, we adopt training rewards to evaluate the training progress. We also test the performance on the test environments to calculate the win rate for SMAC or score rate for GRF.
We compare the performance of our approach with four other baselines. As SHPPO is built based on parameter-shared HAPPO, we take both original HAPPO where each agent has an individual network, and HAPPO (share) where every agent shares the same parameters as baselines. HATRPO [14] is another heterogeneous MARL method, and for a fair comparison, we adopt its parameter-shared version as a baseline. MAPPO [20] is a classic homogeneous MARL method, and in this paper, we also use parameter-shared MAPPO.
Figure 5: Visualization of MMM2 and latent variables. The agents adaptively employ heterogeneous strategies throughout the task. The latent variables are clustered based on these strategies, indicated by circles of different colors: red: stay back to be covered. green: attack in the distance. blue: attract the fire by moving to the front. Each agent is identified by an ID, and their types are distinguished by the shapes in the lower figures: solid circles: Marauders, triangle: Marines, star: Medivac.
TABLE II: Win rate of scalability test on SMAC.
Task* SHPPO (ours) HAPPO** HAPPO (share) MAPPO (share) HATRPO (share)
@@
    When transferring HAPPO, we randomly select a model of the same type for the added agent, or remove extra models when there are fewer units in the new task.
@@
V-B Results on SMAC
We first evaluate SHPPO on the original tasks. The first task is a heterogeneous task MMM2, where each side has three kinds of different units - Marine, Marauder, and Medivac (Fig. 3(a)). They have different health points, hit points, and even special skills. Therefore, heterogeneous strategies will improve the team’s performance by maximizing each agent’s advantages. In Fig. 4(a), HAPPO performs best as it models every agent individually. However, thus HAPPO consumes more parameters and lacks scalability. In contrast, our approach, SHPPO, outperforms all the other parameter-shared baselines without modeling each agent. The training reward and win rate of SHPPO are very close to those of HAPPO, showing that SHPPO has a similar heterogeneous representation ability to HAPPO (See Table I). The performance of the two heterogeneous methods, HAPPO and HATRPO, is the worst when parameters are shared.
We also illustrate the screenshots with the corresponding latent variables (Fig. 5). Surprisingly, the locations of the latent variables can explain the agents’ strategies. At the early stage of the task (Step 5), the agents form two groups, the Marines move forward and attack in the front (green circle) to cover the Marauders and the Medivac behind (red circle). At step 11, the two armies encounter and the strategies alter accordingly. The Marauders get close to the enemies to attract the fire while the Medivac keeps curing them (blue circle). Some Marines (Agent 2, 5, and 8) are hurt with low health points, thus they retreat to be covered (red circle), while the other Marines attack the enemies (green circle). In the final stage, some agents are dead and the enemies are almost defeated. The only Marauder keeps attracting the fire in the front. However, the Marines attack the enemies regardless of their health points to achieve final victory. We can see that the latent variables successfully learn the heterogeneous strategy patterns in this task, which are located diversely in the latent space. These latent variables contribute to the adaptive policy updates during the task through the HeteLayer in the ActorNet, so the agents can not only have distinct strategies but also change their strategies as the task progresses.
Moreover, we evaluate the proposed method on a homogeneous task 8m_vs_9m (Fig. 3(b)), where each side has the same kind of unit - Marine. Though the agents are all the same, they can have different strategies and adaptively alter during the task. As shown in Fig. 4(b) and Table I, SHPPO outperforms all the baselines including HAPPO. The training process of SHPPO also converges faster than other baselines.
@@ V-C Results on GRF
In another task counterattack_easy (Fig. 3(d)), four agents need to cooperate together to do a counterattack. They need to pass the ball against one adversary player and the keeper to score. In Fig. 4(d), SHPPO reaches a similar performance with HAPPO after training, though SHPPO does not explicitly model each agent. Still, SHPPO outperforms all the parameter-shared baselines (also see Table I).
Figure 6: Visualization of the transferred task 821_831 and latent variables. The heterogeneous strategies update accordingly when the population size and task difficulty vary. The notations are consistent with those in Fig. 5.
TABLE III: Score rate of scalability test on GRF.
Task SHPPO (ours) HAPPO HAPPO (share) MAPPO (share) HATRPO (share)
@@ V-D Sacability Test
From Table II, SHPPO performs the best on most of the unseen zero-shot tasks. Even in cases where SHPPO is not optimal, it still performs comparably to the best. In tasks 821_831 and 731_831, there is one more Marine and Marauder, respectively. Therefore, the homogeneity increases after the transfer, while MAPPO is trained with a shared network and may work better when the task is more homogeneous. It’s worth noting that although we attempt to transfer HAPPO, its performance on all the new tasks is worse due to over-fitting to the role assignments in the original task, despite having a better win rate on the original task.
In Fig. 6, we illustrated the latent variables after the transfer (821_831), where one new Marine is added while the enemies are the same, making the new task easier. Thus, the agents inherit the previous strategies and roles in the team, but further adaptively align them with the easier scenario. At step 7, the added Marine (Agent 9) joins its fellow Marines to attack the enemies (green circle), while other Marines are moving forward to attract and disperse the enemies’ fires, together with the Medivac and Marauders (blue circle). Note that no agents keep staying back to be covered as the red circle in Fig. 5, because they are taking a more aggressive strategy in easier tasks. At step 16, all the Marines stop moving and focus on attacks (green circle). The Marauders go on moving to the front with the support of the Medivac (blue circle). At the final stage (step 33), Marauder 1 dies, and almost at the same time, Marine 3 changes its strategy to replace the dead Marauder, moving to the front (blue circle). With the good representation ability of the latent variables, the HeteLayer can inherit the learned strategies in the original task. More importantly, thanks to the adaptive inference of latent variables based on the observations, the agents are able to update collaboration strategies accordingly when the population size and task difficulty vary.
Figure 7: Ablation studies. The experiments are conducted on MMM2 and 8m_vs_9m in SMAC. (a, b) Ablation studies of LatentNet’s inputs. (c, d) Ablation studies of LatentNet’s losses. The confidence interval is calculated over 5 seeds.
@@
Similarly, we test scalability by removing or adding one agent in the two tasks of GRF. SHPPO maintains the best performance after transfer on the 3_vs_1_with_keeper series. For the counterattack_easy series, despite HAPPO’s best performance on the original task, SHPPO achieves the superior score rate on 3_vs_1_counterattack_with_keeper after the transfer. When adding a new agent, SHPPO also has comparable performance.
@@ V-E Ablation Studies
In order to test the effectiveness of the latent variables, we first only mask the inputs of the LatentNet with zeros so that the latent variables cannot be adaptively updated according to the new observations. Moreover, in another experiment, we replace all the latent variables with zeros to totally remove the latent learning. In Fig. 7(a, b), when the inputs of LatentNet are zeros, the performance on two tasks of SMAC reduces significantly, especially in task 8m_vs_9m. When the latent variables are set as zeros, it is equal to the removal of the latent learning part, and we can see even worse performance in such a case. Therefore, flexible adjustment of heterogeneous strategies is important for the agents to cooperate.
@@
Next, we test the effectiveness of the losses of latent learning. We first remove the guidance from the InferenceNet $\mathcal{L}_{v}(\theta_{L})$ while keeping the other losses $\mathcal{L}_{e}(\theta_{L})$ and $\mathcal{L}_{d}(\theta_{L})$. Results in Fig. 7(c, d) show that without $\mathcal{L}_{v}(\theta_{L})$, there will be a performance decline on both MMM2 and 8m_vs_9m. It suggests that the InferenceNet is able to assess the value of latent variables to help with agents’ decision-making.
@@
Similarly, in Fig. 7(c, d), we then keep $\mathcal{L}_{v}(\theta_{L})$ but remove $\mathcal{L}_{e}(\theta_{L})$ and $\mathcal{L}_{d}(\theta_{L})$, respectively. The two experiments have similar performance reductions on task MMM2. Nevertheless, on the homogeneous task 8m_vs_9m, the experiment without the entropy loss $\mathcal{L}_{e}(\theta_{L})$ has the lowest win rate. We believe this is due to the fact that, while strategies for a homogeneous task may exhibit some degree of diversity for labor division, they cannot be excessively random or divergent, as the agents belong to the same unit type. While the entropy loss $\mathcal{L}_{e}(\theta_{L})$ can make sure that the learned latent variables are relatively deterministic and identifiable.